ISS Daily Summary Report – 4/14/2017

Joint Station Local Area Network (LAN) Version 10.0 Transition: This morning the crew replaced the Edge Routers in the Lab and Node 2. Following the hardware replacement, ground teams began intsalling JSL v10 and configuring the network using the upgraded routers. The previously installed Integrated Station LAN (ISL) Routers were reaching end of life on their flash memory and the new upgrades allow support of gigabit ethernet connectivity which will result in increased network throughput. Checkout of the system by ground teams is in work and proceeding nominally.

S11A_C Remote Power Controller (RPC) 3 Status:  Yesterday afternoon, Flight Controllers successfully commanded this RPC closed after a 3rd attempt.  Full activation of the Starboard Thermal Radiator (STR) MDM was then successfully performed. This RPC had initially tripped open on Monday as a result of a Field Effect Transistor (FET) Controller Hybrid (FCH) failure. 

N21B4B_B  RPC 16 Status:  This RPC which powers the Node 2 Moderate Temperature Loop (MTL) System Flow Control Assembly (SFCA) Valve was successfully closed yesterday afternoon. Closure of this RPC enables the ability for closed-loop control of the Node2 MTL loop flow. Ground teams are assessing the root cause of this RPC trip which occurred this past Wednesday.    

Today’s Planned Activities
All activities were completed unless otherwise noted.

  • Node 2 and Lab JSL Router Remove and Replace
  • [СРВК-2М] Gas Liquid Mixture Filter R&R
  • [СРВ-К2М] Filter Reactor R&R
  • Searching for the Causes of Current Overloads in the FGB ПШО31 Exchange Bus Path via [СтА] -Х.
  • Strata Status Check
  • ISSAC Stowage Consolidate
  • CB2 Hardware Relocation
  • ISS HAM Transceiver Swap
  • COTS UHF Communication Unit (CUCU) Activation
  • Manufacturing Device Print Removal, Clean and StowPlug in Plan Hardware Teardown and Stow
  • Photography of Panes of Window #13 in the SM, EV Hatch 1 and EV Hatch 2 in MRM2
  • Plug in Plan Sidekick Stow
  • Biomolecule Sequencer Software Update
  • ISS HAM Service Module Pass
  • Environmental Health System (EHS) Total Organic Carbon Analyzer (TOCA) Water Recovery System (WRS) Sample Analysis
  • CWC-I to EDV Transfer (start)
  • Environmental Health System (EHS) Total Organic Carbon Analyzer (TOCA) Sample Data Record
  • COTS UHF Communication Unit (CUCU) Deactivation
  • CWC-I to EDV Transfer (end)

Completed Task List Items

  • None

Ground Activities
All activities were completed unless otherwise noted.

  • Joint Station LAN Transition
  • JEMRMS SFA Installation 

Three-Day Look Ahead:

  • Saturday, 04/15: Crew off duty, housekeeping
  • Sunday, 04/16: Crew off duty, Marrow
  • Monday, 04/17: OBT 49S Emergency Drill, Cygnus OBT, Neuro Mapping, Meteor hard drive swap

QUICK ISS Status – Environmental Control Group:







[СКВ] 1 – SM Air Conditioner System (“SKV1”)


[СКВ] 2 – SM Air Conditioner System (“SKV2”)


Carbon Dioxide Removal Assembly (CDRA) Lab


Carbon Dioxide Removal Assembly (CDRA) Node 3


Major Constituent Analyzer (MCA) Lab


Major Constituent Analyzer (MCA) Node 3


Oxygen Generation Assembly (OGA)


Urine Processing Assembly (UPA)


Trace Contaminant Control System (TCCS) Lab

Full Up

Trace Contaminant Control System (TCCS) Node 3